OpenStudy (anonymous):

A pier forms an 85° angle with a straight shore. At a distance of 100 feet from the pier, the line of sight to the tip forms a 37° angle. Find the length of the pier to the nearest tenth of a foot.

OpenStudy (amistre64):

100 feet which way? of the pier is at a slant of 85 degrees, then are we 100 feet into the acute or the obtuse angle?

OpenStudy (amistre64):

ill assume the acute angle; and use the law of sines. we are given 2 angles for the triangle and the length of one side (100ft). We find the missing angle by: 180 - a - b = c 180 - 37 - 85 = 58 degrees. the law of sines says the sinA/a = sinB/b = sinC/c angles are in degrees... sin(37) sin(58) ----- = ------- x 100 x = 100 sin(37)/sin(58) the pier is: 70.96 barring any unforseen errors in computations :)